OpenStudy (anonymous):

Which is a solution of the inequality? p + 4 – 2(p – 10) > 0

OpenStudy (eyust707):

Just treat it as if the > was an equals sign

OpenStudy (anonymous):

???

OpenStudy (eyust707):

Do you know how to solve this: \[p+4-2(p-10)= 0\]

OpenStudy (anonymous):

no so i was hoping you can give me the answer haha :)

OpenStudy (eyust707):

Sorry but that violates the overall spirit of this website...

OpenStudy (eyust707):

We want you to learn!! =P

OpenStudy (eyust707):

I spent the time to learn it and use this kind of thing all the time... you should too.

OpenStudy (anonymous):

yes sir >_<

OpenStudy (eyust707):

lol check out what DLS did

OpenStudy (eyust707):

can you follow what he did so far?

OpenStudy (anonymous):

okay is this the answer 26 ??? or 24

OpenStudy (dls):

p + 4 – 2(p – 10) > 0 p + 4 – 2p + 20)> 0 p<24

OpenStudy (dls):

Just treat the ">" symbol as "=" like you solve normal equations.

OpenStudy (anonymous):

so 24 is the answer ???

OpenStudy (anonymous):

and okay got it !!!

OpenStudy (dls):

p + 4 – 2(p – 10) =0 p+4-2p+20=0 You can solve this?

OpenStudy (dls):

Glad :)

OpenStudy (anonymous):

thanks:)